Find the energy of a mosquitoes weighing 0.01g are flying at a speed of 0.5m/s
1
Expert's answer
2019-11-20T09:42:07-0500

The kinetic energy of a mosquito


"KE=\frac{mv^2}{2}"

"=\frac{0.01\times 10^{-3}\:\rm kg\times (0.5\:\rm m/s)^2}{2}=0.00000125\:\rm J=1.25\: \mu J"